Hello! I am trying to solve this one using back substitution.
{x-y-2z=10
{2y+2z=-8
z=-1

Thanks

Answer by Alan3354(69443)

Linear-systems/385954 (2010-12-14 08:47:52): Hello! I am trying to solve this one using back substitution.
x-y-2z=10
2y+2z=-8
z=-1
------------
2y+2z=-8
Sub for z
2y + 2*-1 = -8
2y = -6
y = -3
-------
x-y-2z=10
Sub for y & z
x - (-3) -2(-1) = 10
x + 3 + 2 = 10
x = 5